OpenStudy (anonymous):

having problem with solving given formula A=a+b+c /3 solve for b

OpenStudy (anonymous):

multiply by 3 a•3=a+b+c subtract a 3a-a=b+c 2a=b+c Subtract c 3a-a-c=b